# Jon Danielsson London School of Economics Andreas Uthemann

### v4

@@ 3.2 How do we know AI does what it is supposed to?
One of the hardest problems for AI applied to making decisions in complex social settings, like financial policy, is the specification of its objectives. Why should we trust its analysis and explanations?
A key challenge is that the training algorithms need to know what to optimise for. In other words, they need a well specified objective function that evaluates the cost and benefits of alternative courses of action given the current state of the environment and its future evolution. In most financial applications, they need to take into account how the system reacts to its actions. Misspecifying the problem, or what is likely to be more common, insufficient specification of the problem leads to suboptimal and even catastrophically bad decisions.
In the 1980s, an AI decision support engine called EURISKO used a cute trick to defeat all its human competitors in a naval wargame, sinking its slowest ships to maintain manoeuvrability. This early example of AI reward hacking, something humans are experts in, illustrates how difficult it is to trust AI. How do we know it will do the right thing? Human admirals don’t have to be told not to sink their own ships, and if they do, they either have high-level political acquiescence or are stopped by their junior officers. Any current AI making autonomous decisions has to be told, or learn from observing human decisions, that sinking its own ships is not allowed. The problem is that the real world is far too complex for us to train AI on every eventuality. AI will predictably run into cases where it will make critical decisions in a way that no human would. EURISKO’s creator, Douglas Lenat, notes that “[w]hat EURISKO found were not fundamental rules for fleet and ship design; rather, it uncovered anomalies, fortuitous interactions among rules, unrealistic loopholes that hadn’t been foreseen” (Lenat, 1983, p 82). Each of EURISKO’s three successive victories resulted in rule changes intended to prevent repetition. Still, in the end, the only thing that worked was telling Lenat that his and his AIs’ presence was not welcome.
If we ask ChatGPT whether it is okay for admirals to sink their own ships, it gives a well-grounded, nuanced answer. However, this is a well known example, likely in the canon ChatGPT trained on. Would it have come up with the same answer if Lenat had not entered EURISKO in the naval board game? We don’t know.
@@
AI, which cannot explain its reasoning, cannot be used for many applications. Not surprisingly, considerable resources are being brought to bear on that problem, allowing the operators to trace particular recommendations to the input data, so instead of a recommendation being merely a function of model parameters, they are determined by particular inputs that are analysed logically. Benchmarking AI engines to specific tasks could be particularly valuable in financial applications, as we discuss in Section 6.
4 AI use in the financial system
@@
Many public and private financial system applications can be supported by the two forms of AI accessible today, reactive and limited memory. And, as the engines’ capabilities develop at a rapid pace, and we come closer to a theory of mind, and possibly AGI, we expect AI use to expand in both the public and private sectors (Maple et al., 2023; Kinywamaghana and Steffen, 2021; Korinek, 2023).
Credit allocation, fraud, compliance and risk management are just a few of the tasks in which reactive machines have proven highly valuable. Supervisors may use them to develop the rulebook and enforce compliance with it. There is no technical reason why AI cannot perform the majority of such jobs now; what is holding it back is a lack of human expertise, compute and data sharing (see Section 2.1) as well as caution.
For private financial institutions, where senior managers can be held to account, the unknown territory of AI, compared to relatively safe and familiar incumbent technology, might represent unacceptable personal risk. However, while legacy systems might hold back traditional financial institutions and perhaps a culture that reflects traditional practices, challenger banks find it much easier to use technology. In a competitive market, their greater efficiency and better delivery of financial services will most likely push AI adoption across the private sector.
They may, however, prefer to outsource much AI analytics because of the high fixed costs of AI engines, as discussed in Section 5.4. Outsourcing to a vendor that spreads compute and human capital costs across multiple banks may be especially appealing to all save the largest banks.
The authorities have chosen a more cautious approach to AI. However, as the private sector develops its use of AI, the authorities must follow suit if they are to remain relevant. They are already doing so, as evidenced by Moufakkir (2023) and Cook (2023). Anecdotal evidence suggests that current AI use is mostly limited to regular forecasting and low-level data processing. However, most authorities are only now beginning a more structured analysis of AI.
4.1 Authorities taking advantage of AI
@@
While financial authorities have so far used AI fairly sparingly, we expect that to change. This will necessitate investments in data, human capital and computing power. The most straightforward is data, as authorities already have access to large databases, including confidential disclosures, that can be used to train AI. This includes:
@@
  • Prices, trading volumes, and securities holdings in fixed-income, repo, derivatives and equity markets
@@
    • Central bank speeches, policy decisions, staff analysis
    • Records of past crisis resolution
@@
    • Equilibrium macro model for policy analysis
@@
The authorities will find human resources and compute more difficult, not the least because of very high costs of human capital specialised in AI, as we discuss below in Section 5.4. Training AI engines is very expensive and necessitates access to specialised facilities, both of which are out of reach for most financial authorities. However, because many of the largest costs are associated with training general-purpose engines, it is easy to overstate those problems. The marginal cost to the authorities is likely to be significantly smaller, as they may be able to extend extant engines via transfer learning on curated datasets for financial sector use that include the items listed above.
There are many activities within the financial authorities where AI would be of considerable help. Start with the microprudential regulations, both design and supervision. ML can be used to translate the rulebook into a logic engine, allowing analysis of internal consistency and coverage and facilitating AI supervision. Several authorities have already made progress in this area. Generative models can simulate undesirable behaviour and hence help in designing the rulebook to protect against such outcomes. Reactive engines can identify undesirable behaviour not covered by extant rules and propose remedies.
AI can be beneficial in supervision, providing routine compliance advice to the private sector, recommending regulatory actions, and even making supervisory decisions. While human supervisors would initially closely oversee it, reinforcement learning with human feedback will help supervisory AI to become increasingly performant. For challenging areas of microprudential supervision, adversarial architectures such as generative adversarial networks (GANs) might be particularly beneficial in understanding complex areas of authority-private sector interactions, such as fraud detection.
@@
It will also likely be useful in other activities, such as ordinary economic analysis and forecasting, thanks to an economic engine that performs both theoretical and statistical analysis. This could be achieved with a general purpose foundation model augmented via transfer learning using publicly and privately available data, established economic theory embodied in economic models and previous policy analysis. Reinforcement learning with feedback from human experts might be useful to improve the engine further. Such AI would be very beneficial to those conducting macroprudential stress tests. A key challenge in such analysis (Anderson et al., 2018) is that the stress needs to cover behaviour that we don’t usually see in the system and capture particular interrelations between the various market participants in times of stress. Here, generative models could be valuable in facilitating the running of scenarios in crisis resolution, helping to identify and analyse the drivers of extreme market stress. The authorities could also use such models as artificial labs to experiment on policies and evaluate private sector algorithms, perhaps building on existing agent-based simulations. Over time, as AI technology improves, theory of mind AI might be able to understand and provide context for policy decisions and hence be invaluable in crisis resolution. AI that understands how its interlocutors think can explain its decisions to them, as well as reason about how they will react to its actions and plan accordingly, will be of considerable benefit. Not the least, it could address the problem of misspecified objectives.
4.2 The risk of authorities losing control

### v5

@@ 3.2 How do we know AI does what it is supposed to?
One of the hardest problems for AI applied to advising and making decisions in complex social settings, like financial policy, is the specification of its objectives. Why should we trust its analysis and explanations, and how to make AI systems behave in line with human intentions and values? [^8] For a survey of AI alignment work, see, e.g., Ji et al. (2023), and Shevlane et al. (2023), and Bengio et al. (2023) model evaluation for extreme risks.
A key challenge is that the training algorithms need to know what to optimise for since AI need well-specified objective functions that evaluate the cost and benefits of alternative courses of action given the current state of the environment and its future evolution. That extends to AI taking into account how the system reacts to its actions, which is particularly important in financial applications. Misspecification, and especially insufficient specification of problems, leads to suboptimal and even catastrophically bad decisions.
In the 1980s, an AI decision support engine called EURISKO [^9] Following the passing of EURISKO’s creator, Douglas Lenat, the source code was re-discovered and put online, blog.funcall.org/lisp/2024/03/22/eurisko-lives used a cute trick to defeat all its human competitors in a naval wargame, sinking its slowest ships to maintain manoeuvrability. This early example of AI reward hacking, something humans are experts in, illustrates how difficult it is to trust AI. How do we know it will do the right thing? Human admirals don’t have to be told not to sink their own ships, and if they do, they either have high-level political acquiescence or are stopped by their junior officers. Any current AI making autonomous decisions has to be told, or learn from observing human decisions, that sinking its own ships is not allowed. The problem is that the real world is far too complex for us to train AI on every eventuality. AI will predictably run into cases where it will make critical decisions in a way that no human would. EURISKO’s creator, Douglas Lenat, notes that “[w]hat EURISKO found were not fundamental rules for fleet and ship design; rather, it uncovered anomalies, fortuitous interactions among rules, unrealistic loopholes that hadn’t been foreseen” (Lenat, 1983, p 82). Each of EURISKO’s three successive victories resulted in rule changes intended to prevent repetition. Still, in the end, the only thing that worked was telling Lenat that his and his AIs’ presence was not welcome.
If we ask ChatGPT whether it is okay for admirals to sink their own ships, it gives a well-grounded and nuanced answer. However, this is a well known example, likely in the canon ChatGPT trained on. Would it have come up with the same answer if Lenat had not entered EURISKO in the naval board game? We don’t know.
@@
AI, which cannot explain its reasoning, can only be used for some applications. Not surprisingly, considerable resources are being brought to bear on that problem, allowing the operators to trace particular recommendations to the input data. Instead of a recommendation being merely a function of model parameters, recommendations are determined by particular inputs that are analysed logically. Benchmarking AI engines to specific tasks could be particularly valuable in financial applications, as we discuss below.
4 The channels for how AI can destabilise the financial system
The reason why AI increases systemic risk is, perhaps paradoxically, its very strength and efficiency. In the words of Bengio et al. (2023), “Compared to humans, AI systems can act faster, absorb more knowledge, and communicate at a far higher bandwidth. Additionally, they can be scaled to use immense computational resources and can be replicated by the millions.” The financial system has a large number of potentially destabilising feedback mechanisms that are created by the interaction of actors with significant financial resources that are maximising profits in regular time and the probability of survival during stress. The combination of AI that excels at discovering optimal strategies, the system’s complexity, highly resourced strategic actors and the inherent destabilising factors already prevalent in the financial system viciously reinforce each other.
In our analysis of how AI can destabilise the financial system, we build on extant work on AI safety, like Weidinger et al. (2022), Bengio et al. (2023) and Shevlane et al. (2023), which identifies several societal risks arising from AI use, including malicious use, misinformation and loss of human control. In addition, AI brings a particular market structure. When adapting these AI risks to the financial sector and analysing them in the context of the financial vulnerabilities developed above in Section 2, we arrive at four channels for how AI can destabilise the financial system.
4.1 Malicious use of AI
The first channel arises because the system is replete with highly resourced profit-maximising economic agents not too concerned about the social consequences of their activities. They seek to bypass controls and change the system in ways that benefit them while being difficult for other market participants and regulators to detect. Technology has always helped such agents, and they are already embracing AI. Financial institutions and their staff deploy considerable resources in their pursuit of profit and usually are not very concerned about the wider social impact of their activities. They can change the system in ways that benefit them while not being detectable by others, as we have seen many times in the past. AI provides such actors with new opportunities, perhaps via adversarial attacks by feeding particular data into others’ training algorithms or influencing market structure. That might not even be necessary since it is straightforward to exploit for private, even illicit, gain gaps in how AI sees its responsibilities where “maximise profit” leads to market manipulation without explicit instruction by humans.
One way such malicious use might occur is by manipulating governance processes. The financial authorities and internal control aim to align the interests of system participants with those of the organisation that employs them and society. As AI is particularly effective in finding profitable loopholes and amplifying vulnerabilities, it can facilitate misbehaviour that, while legal, is damaging both to society and even the institution employing the AI.
The operator of the AI could take that one step further and use it to deliberately create market stress, which, of course, is profitable to those forewarned. We have seen many non-AI examples of such conduct in the past, and a key purpose of securities and macroprudential regulations is to prevent such behaviour.
We expect the most common malicious use of AI will be by employees of financial institutions, who are careful to stay on the right side of the law. AI will also facilitate illegal activities, such as rogue traders and criminals, as well as terrorists and nation-states aiming to create social disorder. The financial system has always attracted such hostile agents, and nation-states view financial system attacks as one part of their arsenal, as we have witnessed many times in the past. The inherent vulnerability of the financial system to adversarial attacks and the ease of manipulating AI engines create potential for particular attack vectors that can be impossible to identify ex-ante. One example comes from Hubinger et al. (2024), who identify what they call a “sleeper agent” that usually acts as expected but gives deceptive answers when given special instructions. Here, AI engine corruption becomes another form of cyber attack. With such an attack vector in place, an AI engine becomes a tool for well-resourced hostile agents that might be difficult or impossible to prevent in real-time, facilitating nation-state attacks on critical infrastructure. That is particularly relevant when a superior ability to identify optimal timing strategies allows them to solve the problem of double coincidence, as discussed in Daníelsson et al. (2016), creating heightened system fragility, perhaps by manufacturing a liquidity crisis as an attack amplifier.
4.2 Misinformed use and overreliance on AI
The second channel emerges when the human operators of AI engines don’t fully appreciate how AI operates and then end up using it inappropriately. As we come to trust AI analysis and decisions and see how well it performs in increasingly complex and essential tasks, it builds up trust. There are many examples where technology was initially met with scepticism but then became trusted, with just some recent examples: computers landing aeroplanes, doing surgery and driving cars. While advice from AI might be reviewed by humans, once we see it suggest the right thing many times, eventually, it becomes accepted. When AI performs at a level equal to or better than its human counterparts but at a much lower cost, it will be welcomed by senior management. Its very success creates trust. As that trust is earned on relatively simple and safe repetitive tasks and seeing AI succeed in more complex jobs, it will be asked to take on ever more sophisticated tasks.
That is particularly relevant when AI is used in domains with unclear objectives, which is very common in the financial system. As events become less frequent and more serious, data-driven decisions increasingly involve extrapolation, interpretation, intuition and nuance, none of which AI in its current form is good at. As we discuss on Page 2.4, many crises have been resolved by very creative solutions that had never been used before.
Because successful AI architectures are data-driven engines, their analysis is predominantly founded on statistical regularities. Such AI engines are designed to provide advice and responses to prompts, even if they have very low confidence about the accuracy of the answer. Of course, many architectures are probabilistic and can provide measures of statistical accuracy. They can even make up facts or present arguments that sound plausible but would be considered flawed by an expert, both instances of the broader phenomenon of AI hallucination. The risk is that the AI engines will present confident recommendations about outcomes they know little about.
4.3 AI misalignment and control avoidance
The problem of misalignment is particularly difficult when several AI engines interact. For example, market manipulation is most successful when several investors collude. Similarly, the payoff of many trading activities, such as carry trades and short sales, increases with the number of investors participating in them. A particularly damaging consequence of the strategic complementarities prevalent in the financial system relates to the objectives of financial institutions. They maximise profit most of the time, perhaps 999 days out of a thousand, but optimise for survival during stress. The behaviour on that one day out of a thousand is particularly damaging when financial institutions seek safety by withdrawing liquidity, which can lead to destabilising dynamics such as fire sales, bank runs, and credit crunches. The very high level of AI performance can, perhaps paradoxically, increase the likelihood of damaging coordinated behaviour. They might find better strategies for exploiting strategic complementarities for private gain at the risk of lowering market quality or even threatening financial stability.
An example of AI collusion is Calvano et al. (2020), who find that independent reinforcement learning algorithms instructed to maximise profits quickly converge on collusive pricing strategies that sustain anti-competitive outcomes. It is much easier for AI to behave in this collusive way than humans, as such behaviour is very complex and often illegal. AI is much better at handling complexity and is only aware of the legal nuances if explicitly taught or instructed.
The most difficult problems of misalignment arise due to the interaction of the various AI working throughout the industry. Long before computers got involved with the financial system, humans coordinated damaging behaviour due to the misalignment of individual short-term incentives with the long-term incentives of the same market participants and society at large. The crises of 1763, 1914 and 1929 (see, e.g. Daníelsson, 2022) are just two of many such examples, as is the tulip mania in the Netherlands in the 1630s, see Aliber and Kindleberger (2015).
The use of computers amplifies the potential for such undesirable outcomes, as two recent examples illustrate. The largest stock market crash in history happened in 1987 when global markets fell by 23% in one day. The reason was trading algorithms that, in a way not foreseen, coordinated in creating vicious feedback between prices collapsing and algorithm-induced selling, as documented by Gennotte and Leland (1990). Similarly, the quant crisis in the summer of 2007, as noted by Khandani and Lo (2007), was due to trading algorithms trained to sell on falling prices but not being told of the potential for aggregate feedback caused by trading algorithms in a number of funds coordinating on selling. Such factors have further been at the root of the various flash crashes. AI will likely amplify such problems of misalignment. AI might implement cross-market strategies that no humans will be able to understand, let alone unwind. While the traditional approach to algo (or human) induced market crashes is trading stops, i.e., a simple kill switch, that might not be feasible for complex cross-market AI strategies.
4.4 Market structure
Monoculture is an important driver of booms and busts in the financial system. As financial institutions come to see and react to the world in increasingly similar ways, the more they coordinate in buying and selling, leading to bubbles and crashes or, even worse, credit booms and credit crunches. aucommFor me, credit booms and crunches can be instances of bubbles and crashes. Banks’ capital calculation is a particular example, not the least via risk weights. Such procyclical behaviour has always been a feature of financial markets, long before the use of computer technology. Still, it is amplified by technology.
Finally, financial data vendors have concentrated considerably over the past few years, with only a few large vendors left, such as S&P Global, Bloomberg and LSEG. They have unified databases with standardised labelling and graph structures linking disparate categories of data, including proprietary and not publicly available data, facilitating engine training. It is a concern that neither the competition nor the financial authorities have fully appreciated the potential for increased systemic risk due to oligopolistic AI technology in the recent wave of data vendor mergers.
Taken together, AI-driven financial analytics is characterised by increasing returns to scale technologies with high entry costs, making an oligopolistic market structure highly likely. In such an environment, outsourcing analytics to one of those vendors might be the best, maybe even the only feasible solution for all but the largest financial companies. Anecdotal private sector evidence indicates that many financial institutions already prefer to outsource data-driven applications to specialists. A particular example is risk management as a service, RMaaS, with a leading vendor being BlackRock’s AI-fuelled Aladdin.
The risk is harmonised beliefs and actions, driving procyclical behaviour and exposing users to the same blind spots.
5 The use and regulation of AI
The financial system has always been an early and enthusiastic adopter of technology. One of the first applications of the first transatlantic telegram cable in 1858 was the transmission of stock prices. Nathan Rothschild supposedly used pigeons to get the first news of Napoleon’s defeat at Waterloo in 1815 in order to manipulate the London stock market. As many such examples show, technology is both a source of efficiency and risk, and we now see the same with AI.
@@ 5 The use and regulation of AI
Most public and private financial system applications can be supported by the two forms of AI accessible today, reactive and limited memory, and their associated architectures. As the engines’ capabilities improve, and we come closer to a theory of mind AI, and even AGI, we expect AI use to expand in both the public and private sectors (Maple et al., 2023; Kinywamaghana and Steffen, 2021; Korinek, 2023).
Most job functions in the public and private parts of the financial system are quite routine, and there is no technical reason why AI today cannot perform the majority of these jobs. Credit allocation, customer interactions, report generation, pricing of insurance contracts, fraud detection, compliance and risk management are just a few of the tasks where existing AI have proven their worth. While many financial institutions are making large investments in AI, what often holds AI adoption back is lack of familiarity with it and insufficient access to human resources, data and compute, and perhaps most importantly, caution based on worries about what sort of mistakes AI advice and decisions could lead to. Many financial institutions lack the necessary resources, and smaller legacy institutions can find AI particularly challenging. Legacy systems and processes might not be well suited to AI adoption.
These issues don’t generally apply nearly as much to challenger financial institutions, such as the various neo-banks. These are new institutions with modern technology stacks, benefitting from technology staff and senior management that are conversant with modern technology. In a competitive market, greater efficiency and better delivery of financial services due to AI will push its adoption across the private sector.
Since the financial system is one of the most heavily regulated parts of the economy, the financial authorities similarly take a keen interest in how technology is used. They want to identify the threats technology poses, how to harness it to improve the system and how to regulate it. A particular challenge for the architects of financial regulations is that they simultaneously need to keep the system safe and make it more efficient. As the private sector increasingly adopts AI, the supervisors will have to address both how it changes the behaviour of private sector institutions and how to regulate AI engines in use.
Regulating the financial system is not an easy task. The authorities have to balance conflicting objectives while wielding limited resources when patrolling a nearly infinitely complex financial system. Once the authorities are focused on particular activities, the resulting risk might be contained, but what tends to happen is that the forces of instability emerge elsewhere. Risk is like a balloon; squeeze it in one area, and it expands elsewhere. Indeed, it is almost axiomatic that instability arises where the authorities are not looking. And there is plenty of scope for instability in an infinitely complex system. The authorities will have to contend with how AI changes regulations, how to regulate AI and perhaps most importantly, how to harness it in their operations.
5.1 The authorities’ use of AI
Most financial authorities are approaching AI cautiously and are behind the private sector in both analysing how AI changes the system and how to use it in their activities. Anecdotal evidence suggests that their current AI use is mostly limited to routine forecasting, basic analysis with LLMs and low-level data processing. Some authorities are increasingly considering how AI will affect their mission, as evidenced by Leitner et al. (2024), Moufakkir (2023) and Cook (2023). If they don’t keep up, the authorities risk becoming increasingly irrelevant if their way of working, such when monitoring, regulating and supervising the financial system, is not geared towards private sector use of AI. The financial authorities find AI challenging. Engaging with it necessitates investments in data, human capital and compute, as well as demanding re-thinking in how to regulate finance.
@@ 5.1 The authorities’ use of AI
The most straightforward is data, as authorities already have access to large databases, models and technical processes, including confidential disclosures, that can be used to train AI. This includes:
  1. 1.
@@
  2. 2.
    Prices, trading volumes and securities holdings in fixed-income, repo, derivatives and equity markets
  3. 3.
@@
    2. (b)
      Speeches, policy decisions, staff analysis
    3. (c)
      Records of past crises resolution
  7. 7.
@@
    4. (d)
      Equilibrium macro models
    5. (e)
      System stability models
@@
By contrast, the authorities can find it difficult to acquire the necessary human resources and compute, not the least because of the very high costs we discuss above in Section 4.4. However, because many of the largest costs are associated with training general-purpose engines, it is easy to overstate those problems. The marginal cost to the authorities is likely to be significantly smaller. To begin with, they may be able to extend extant engines via transfer learning on curated datasets for financial sector use that include the items listed above. Furthermore, open-source platforms running on relatively low-end computers are sufficient for many applications.
There are many activities within the financial authorities where AI would be of considerable help. Start with the design of microprudential regulations. ML can be used to translate the rulebook into a logic engine, allowing analysis of internal consistency and coverage and facilitating AI supervision, and then enforcing compliance with it. Several authorities have already made progress in this area. Undesirable behaviour, including that not covered by extant rules, can be simulated in order to make the rulebook robust against such outcomes. AI can help in supervision, providing routine compliance advice to the private sector, recommending regulatory actions, and even making supervisory decisions. While human supervisors would initially closely oversee it, reinforcement learning with human feedback will help supervisory AI to become increasingly performant, and repeated success will build trust. This could be useful in resolution, such as finding the least-cost resolution by evaluating a number of alternative resolution approaches. For challenging areas of microprudential supervision, some types of AI might be particularly beneficial in understanding complex areas of authority-private sector interactions, such as fraud detection and money laundering. Hostile agents make extensive use of AI to hide their activities and find new areas to exploit. The authorities, along with private sector advisory firms, are actively counteracting that with AI.
AI will also be useful in other routine activities, such as forecasting and ordinary economic and system fragility analysis. That may use public and private data, established economic theory embodied in economic models and previous policy analysis. Of particular interest when AI is to be used for economic analysis relates to the Lucas (1976) critique, which states that the decision rules used by economic agents depend on the underlying economic environment and can change as regulations change, undermining their effectiveness. Prior to the 1970s, economists believed in large general models informed by statistical patterns, with the Phillips curve, representing a trade-off between unemployment and inflation based on a negative empirical correlation of these two variables, as a centrepiece. [^12] For an important example, see Rasche and Shapiro (1968) explaining the MIT-Fed model used by the Federal Reserve and developed in collaboration with MIT. The contribution of Lucas was to recognise that when the authorities try to exploit this apparent trade-off, they change the way the private sector negotiates wages and sets prices, leading to a breakdown of previously identified correlations and assumed causalities. Attempts to stimulate the economy through this channel create inflation without reducing unemployment, giving rise to the word stagflation.
In response to the Lucas critique, the economic profession moved away from large-scale simultaneous equation models based on statistical correlations to smaller structural models that explicitly model beliefs and general equilibrium effects so that reactions to policy interventions follow a well-understood and internally consistent logic encoded in the model. A key benefit is that such models provide explainability. AI engines might have to do the same if they are to become effective economic analysts and certainly macropru regulators.
That leaves the question of whether AI will help in creating more flexible large statistical models based on neural networks. That, in part, depends on whether the appropriate policy experiments are in its training dataset or not. If so, the AI engine can identify the various reaction functions, overcoming the Lucas critique. Since a large number of policy experiments have been conducted, an AI model should significantly outperform the old-style statistical models. That, however, does not apply to experiments not in the training dataset, following from Sims’ (1980) comparison of structural and reduced-form models. He notes that this problem can be partially overcome by using reduced-form models for policy analysis if the data contains policy changes. Interpreting his suggestion for the AI era, AI will significantly expand the domain of data-driven policy analysis as long as we recognise the boundaries for sound use and ensure policy experiments are in its training dataset.
In macropru applications, such as stress testing, reinforcement learning with feedback from human experts might be useful to improve the engine further. A key challenge in such analysis (Anderson et al., 2018) is that the stress needs to cover behaviour that we don’t usually see in the system and capture particular interrelations between the various market participants that only emerge in times of stress. Here, some AI models could be valuable in facilitating the running of scenarios in crisis resolution, helping to identify and analyse the drivers of extreme market stress.
@@
The authorities could also use AI models as artificial labs to experiment on policies and evaluate private sector algorithms, building on existing agent-based simulations. Over time, as AI technology improves, theory of mind AI can understand and provide context for policy decisions and hence be invaluable in crisis resolution. AI that understands how its interlocutors think can explain its decisions to them, as well as reason about how they will react to its actions and plan accordingly, will be of considerable benefit. Not the least, it could address the problem of misspecified objectives.
5.2 The authorities’ challenges
While authorities will find AI indispensable in keeping the system stable, they also have to contend with it aiding the forces of instability. Which of these two dominates? Most likely instability, as it needs only to find one weakness, while the authorities must monitor the entire system. The supervisor must not only identify the pressure points but also monitor how financial institutions interact with each weakness in real-time. It is a much more difficult computational task, and one where the more complex the system becomes, the harder the task is. Furthermore, the supervisors have fewer resources than the private sector.
AI creates new forms of model risk, defined as the potential for undesirable outcomes from inaccurate or misapplied models. In the words of George Box (1976), “Every model is wrong, some are useful,” and the problem arises because we are trying to map a problem that is effectively infinitely complex onto a finite representation of the world. Model risk is present in all financial models and is exacerbated by AI because of its complexity and opacity, which makes neural networks particularly difficult to interpret, so it becomes harder to detect errors. The complexity of AI models demands particularly robust validation and governance frameworks. These problems can be mitigated by explainability, validation, and monitoring, as well as by using representative data and validating analysis by alternative architectures. It will be important to have a clear view of type 1 and type 2 errors in order to aid the interpretability of AI recommendations. Engines should be required to provide an assessment of the statistical accuracy of their recommendations. Here, it will be helpful if the authorities overcome their frequent reluctance to adopt consistent quantitative frameworks for measuring and reporting on the statistical accuracy of their data-based inputs and outputs.
Both the public and private sectors will likely end up outsourcing analytics to the same handful of very large AI vendors. This blurs the regulator/regulated divide since if the private sector is acting on the same advice as the public sector, the nature of supervision will change. While already an issue with a small number of vendors handling credit risk, reporting, compliance, fraud and the like, the oligopoly of AI vendors creates particular problems. The reason is that a vendor’s engine will likely have a single representation of the financial system, that is, of the stochastic processes that drive market outcomes. This means that advice provided to all users of the engine, in the private and public sectors, will share the same underlying view of the system. That has two consequences. The first is that it will harmonise beliefs, which is pro-cyclical. Furthermore, since the engine is just a model that maps a near infinitely complex system to a simplified version of it, if a private firm unknowingly engages in undesirable actions at the recommendation of the engine, the authority using the same engine might not identify that misconduct. The use of a single engine across a broad range of supervisory activities means mistakes become systemic, not idiosyncratic. To guard against both issues, the authorities should pay particular attention to problems of model diversity and robustness, aim to use different engines than regulated entities and take advice from several vendors on important questions of system stability.
Increased use of private sector AI can result in a new layer of deniability. Assigning legal responsibility for misconduct in the financial sector is already tricky. It will be harder when AI is used to make decisions. Suppose a human operator deliberately instructs AI to break the law for criminal or terrorist purposes or just turns a blind eye to the AI doing so as a byproduct of maximising profits. Even if detected, it might be easily explainable as an unintended and unexpected innocent behaviour. For example, while a rogue human trader can be prosecuted, we cannot do that with a rogue AI trader. The institution employing such AI might profit significantly. Consequently, the increased use of AI in the private sector facilitates the job of those seeking to utilise AI for nefarious purposes by providing them with yet another level of denial, at least until the law and regulations catch up.
A particularly difficult problem relates to how AI changes the way the authorities view their control of the system. The more sophisticated AI engines become, the more difficult it will be to exercise authority over them. After all, the controllers need to understand the AI objectives to design the correct incentives. Given the current state of interpretability of foundation models, this is challenging.